OpenStudy (anonymous):

A cube has a surface area of 6x squared+48x+96 find the length of one side

hero (hero):

Basically you have to factor it

OpenStudy (saifoo.khan):

\[6 (x+4)(x+4)\] side = (x+4)

hero (hero):

I stand corrected. Formula for area of cube is 6s^2 good job

hero (hero):

But you should have wrote it as 6(x+4)^2

OpenStudy (saifoo.khan):

same thing.

hero (hero):

Not really. The form I put it in preserves the form 6s^2

hero (hero):

where s = (x+4)

OpenStudy (anonymous):

answer was x+4

hero (hero):

We know

hero (hero):

That's what s = (x+4) means

hero (hero):

s meaning "side"

hero (hero):

By the way, I wasn't kidding about that pic you have Saifoo. It doesn't suit you. I'm saying that as a friend, not an enemy.

OpenStudy (saifoo.khan):

LOL, alright.. i will bring that into consideration.

hero (hero):

Let me guess... It's an animation of your favorite cartoon character.

OpenStudy (saifoo.khan):

That is not my "favorite" cartoon. He's Dexter. never heard pf them?

OpenStudy (saifoo.khan):

of*

hero (hero):

I could care less who he is. In my world, Dexter doesn't exist.

OpenStudy (saifoo.khan):

good for u.

hero (hero):

I hope I didn't offend you by that

OpenStudy (saifoo.khan):

no u didn't.